OpenStudy (anonymous):

GIVING MEDALS!!!! Find the value of y so that the line passing through (9,0) and (3,y) Has a slope of -1/2 :D

OpenStudy (solomonzelman):

y=-3

OpenStudy (anonymous):

\[\left( \frac{ y2-y1 }{ x2-x1 } \right)= -1/2\]

OpenStudy (solomonzelman):